Javits & Sons' common stock currently trades at $33.00 a share. It is expected to pay an annual dividend of $1.25 a share at the end of the year (D1 = $1.25), and the constant growth rate is 4% a year.

A) What is the company's cost of common equity if all of its equity comes from retained earnings? Round your answer to two decimal places. %

B) If the company were to issue new stock, it would incur a 12% flotation cost. What would the cost of equity from new stock be? Round your answer to two decimal places. %

The Patrick Company's year-end balance sheet is shown below. Its cost of common equity is 17%, its before-tax cost of debt is 12%, and its marginal tax rate is 40%. Assume that the firm's long-term debt sells at par value. The firm’s total debt, which is the sum of the company’s short-term debt and long-term debt, equals $1,191. The firm has 576 shares of common stock outstanding that sell for $4.00 per share. Calculate Patrick's WACC using market value weights. Round your answer to two decimal places. %

Assets Liabilities And Equity
Cash $ 120 Accounts payable and accruals $ 10
Accounts receivable 240 Short-term debt 51
Inventories 360 Long-term debt $1,140
Plant and equipment, net 2,160 Common equity 1,679
Total assets $2,880 Total liabilities and equity

$2,880

Adams Corporation is considering four average-risk projects with the following costs and rates of return:

Project Cost Expected Rate of Return
1 $2,000 16.00%
2 3,000 15.00
3 5,000 13.75
4 2,000 12.50

The company estimates that it can issue debt at a rate of rd = 10%, and its tax rate is 35%. It can issue preferred stock that pays a constant dividend of $3 per year at $53 per share. Also, its common stock currently sells for $33 per share; the next expected dividend, D1, is $3.50; and the dividend is expected to grow at a constant rate of 6% per year. The target capital structure consists of 75% common stock, 15% debt, and 10% preferred stock.

  1. What is the cost of each of the capital components? Round your answers to two decimal places.
    Cost of debt %
    Cost of preferred stock %
    Cost of retained earnings %
  2. What is Adams' WACC? Round your answer to two decimal places.
    %
  3. Only projects with expected returns that exceed WACC will be accepted. Which projects should Adams accept?
    Project 1 -Select-acceptreject
    Project 2 -Select-acceptreject
    Project 3 -Select-acceptreject
    Project 4 -Select-acceptreject
    Answer #1


    Solution: 01 rs=(D1/PO)+g rs=(1.25/33)+0.04 rs= 7.79% PO= market value per share rs = cost of internal equity g= growth rate Q2 Financing Debt Equity Market value Market weight after tax cost of capital weighted cost b=a/3495 d=b*c 1191 0.34077 7.20%Q3 after tax cost of debt (rd) =0.1*(1-0.35) 6.50% cost of preferred stock = rp= Dividend/current price = 3/53 | 5.66% cost o